OpenStudy (anonymous):

Find the first, fourth, and tenth terms of the arithmetic sequence described by the given rule. A(n) = 1 + (n – 1)(-5.7) A) 1, -16.1, -50.3 B) -5.7, -21.8, -51.3 C) 1, -21.8, -56 D) 0, -17.1, -51.3

OpenStudy (eyust707):

plug in a 1 for n

OpenStudy (eyust707):

plug in a 4 for n

OpenStudy (eyust707):

plug in a 10 for n

OpenStudy (anonymous):

A(1) = 1 + (1 - 1)(-5.7) = 1 + (0)(-5.7) = 1 + 0 = 1

OpenStudy (anonymous):

A(4) = 1 + (4 -1)(-5.7) = 1 + (3)(-5.7) = 1 + 17.1 = 18.1

OpenStudy (anonymous):

@eyust707

OpenStudy (eyust707):

Are you sure its written correctly?

OpenStudy (anonymous):

I think so...... What did I do incorrectly?

OpenStudy (eyust707):

youre math is fine thats why im wondering if maybe you copied the problem down wrong

OpenStudy (anonymous):

Well can you help me out?

OpenStudy (eyust707):

Can you take a screenshot or picture of the actual problem

OpenStudy (anonymous):

I don't know how?

OpenStudy (eyust707):

is it online homework?